OpenStudy (anonymous):

Factor completely: 3x3 – 48x A. x(x – 4)(x – 4) B. 3(x + 4)(x – 4) C. 3x(x – 16) D. 3x(x + 4)(x – 4)

Mehek (mehek14):

what do you think?

OpenStudy (anonymous):

i'm thinking either C or D

OpenStudy (michele_laino):

hint: we can factor out 3x: \[\Large 3{x^3} - 48x = 3x\left( {...} \right)\]

Mehek (mehek14):

you have to factor -16 as well

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

or you know no not answer my question

Mehek (mehek14):

what?

OpenStudy (michele_laino):

what is: \[\frac{{3{x^3}}}{{3x}} = ...?\]